Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP:; phone: +1 575-835-7370


News, Travel, Visitors (McMullin)

  • Overall status: B configuration; 4-band dipoles deployed. Correlator hw resources support up to 64 BlBprs (that many sub-bands or use for additional resolution). Stable software system; move begins next week. 11B cycle disposition letters 31 May; next call for proposals 1 Jul.
    • PEPs (self-evaluation) due May 2nd.
  • Key dates:
    • 07 May - 15 Jun: 4-band science observing
    • 09 May: Planned end of B configuration
    • 13-31 May: BnA configuration
  • Staff Travel:
  • Visitors:
    • Robert Harris (AC982): 24 Apr - 16 Jul
    • Alison Peck (ALMA): 15 May - 17 May
    • Bob Sault: 16 May - 23 Jun

Hardware status and updates (Langley, Perley, Rupen)

  • Array (Perley/Langley):
    • EA18 out until 23 May.
  • WIDAR (Rupen):
    • Notes (Rupen/Sowinski): Ongoing focus on 3-bit samplers (report Langley/Dhawan); able to do 64 BlB pairs (two boards need repair/replacement but not affecting data).
  • Computing (Robnett):

Software status and updates (Butler)

  • Links:
  • CBE looking good; pushing ahead on subarrays; CM work to fix an issue recovering from stress/pointing files; CRM delayed to this week (pending adequate testing time). OPT 1.09 testing under webtest is ongoing.
  • CASA Update (new functionality available/issues/enhancements):Notes (Miriam): Stable deployed last Friday; release pending.
  • AIPS Update (new functionality available/issues/enhancements): Notes (Eric):
  • SDM Update: Notes: No news.

Commissioning Group Reports and Test Planning

  • NOTE: Please cc me on submitted tests so that it can be assigned the appropriate priority!

Commissioning priorities for the week (McMullin, Chandler)

Single maintenance day week:
  • Mon
    • 1100-1700 Testing
    • 1800-0900(Tues) Science operations and testing
  • Tue:
    • 0900-1700 Testing/software development time
    • 1800-0630: Science operations and testing
  • Wed: Maintenance day
    • 0630-1700: Maintenance operations
    • 1800-0900: Science operations and testing
  • Thu:
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Fri (holiday):
    • Science observing and testing
  • Sat-Sun
    • Science operations + planned testing

Issues Affecting Science Observing:

  • Correlator: [Pokorny/Rupen/McMullin] time mislabeling in SDM data (non-synchronous operation between correlator dumps and CBE can require additional 1/2 integration time padding for online flags).
  • Correlator: [Clark/Carlson] Delay rocking across bands. Implementation pending factor-of-2 reconciliation.
  • Array: [Gustaaf] Amplitude wobbles seen on subset of antennas; phase jumps at scan boundaries also seen on subset of antennas (Ka, Q bands only thus far).
  • Array: [All] Zoo of other effects (some data are very good, some still seeing antenna-based IF/Polarization issue).

Testing This week

  • 3-bit samplers (Dhawan+): statistics and noise testing (highest priority)
  • Spectral Line (Linz/Ott): Spectral dynamic range test (TSPE0010.3923557). Submitted.
  • Spectral Line (Goss/Chomiuk): BP stability/repeatability (TSPE0005.3699117). Submitted..Completed. Report pending.
  • Correlator: Rupen: S-band shadow testing; Delayed until Michael's return.
  • Correlator: *; L-band SEFD confirmation
    • We present sensitivity measurements of the EVLA at L-band between 0.95 and 2 GHz. These are obtained using three EVLA antennas that are equipped with the final, EVLA-compliant L-band receivers. The reported SEFD values meet and/or exceed the EVLA sensitivity requirements over about 70% of the band, but are too high by 12-38% over the frequency range 1425-1750 MHz.
  • Correlator: Durand/McMullin/Pokorny/Butler; review of CRC error events (may involve testing).
  • Spectral line: Momjian: HI High resolution bandpass testing
  • Array: *; Individual issue sleuthing

Commissioning Groups

    • Adv: Astrometry (Mioduszewski):
      • Notes:
    • Adv: Low band (Owen):
      • Notes: Status report and readiness.
  • Calibrator Models
    • Notes: Status report and Ka/S update.
  • HPC (Bhatnagar/Owen/Golap):
    • Notes:
  • Imaging (Ott/Golap):
    • Notes:


RSRO B configuration status:

Individual reports (updates)


Demo Science Status


  • OPT enhancements for 1.09
    • EVL-1622: Update X band frequencies to full EVLA range
    • EVL-1624: Enable maximum elevation warning (>75 degrees) for Scheduling Block Summary report (for assumed scheduled start time).
    • EVL-1625: Add maximum elevation check to validation (cycling over LST start range)
    • EVL-1626: Enable specification of environmental constraints for the sun (minimum separation, avoid sunrise, avoid sunset)
    • EVL-1627: Enable notification of sunrise or sunset during SB summary report
    • EVL-1620: Incorporate knowledge of known RFI-producing satellites into warning/constraint system (GPS, Iridium, etc)
    • EVL-1621: Add default setups for 4-band

Misc Actions

  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley/McMullin]: SPW/IF/BB lexicon and mapper
  • Development of WIDAR guidelines document (Michael)
  • List of known problems (can be used by OSRO commissioning team (van Moorsel)
      • Med Priority [TBD]: Added. Phase change when frequency changes (needs to be tested).
-- JosephMcMullin - 2011-05-02
Topic revision: r2 - 2011-05-03, JosephMcMullin
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback